OpenStudy (anonymous):

simplify 64^(-1/3)power -4 -1/4 1/4 4 i will reward medal

OpenStudy (anonymous):

@shrutipande9

OpenStudy (shrutipande9):

idk...:(

OpenStudy (anonymous):

it is -one third power

jigglypuff314 (jigglypuff314):

for example: 8^(-1/3) = 1/(8^(1/3)) = 1/2 27^(-1/3) = 1/(27^(1/3)) = 1/3 so 64^(-1/3) = 1/(64^1/3)) = 2*2*2 = 8 8^(1/3) = 2 3*3*3 = 27 27^(1/3) = 3 4*4*4 = 64 64^(1/3) =

OpenStudy (anonymous):

can you explain how to solve step by step it for me cause i want to try and solve it

jigglypuff314 (jigglypuff314):

\[64^{-1/3} \rightarrow \frac{ 1 }{ 64^{1/3} }\]

OpenStudy (anonymous):

ok then what would i do

jigglypuff314 (jigglypuff314):

just like how 16^(1/2) = 4 because 4*4 = 16 8^(1/3) = 2 because 2*2*2 = 8

OpenStudy (anonymous):

ok so i divide 64 by 3?

jigglypuff314 (jigglypuff314):

no... it is to the power so find a number multiplied three times with itself like 4*4*4 = 64

OpenStudy (anonymous):

i got 4 when i cuberooted it

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

so the answer would be 1/4?

jigglypuff314 (jigglypuff314):

yes :) :D

OpenStudy (anonymous):

ok thank you man

jigglypuff314 (jigglypuff314):

your welcome :)